# A Motivic Pairing and the Mellin Transform in Function Fields

### v2

@@ 5. Product Formulas for Exp and Log
In this section we prove formulas which express the exponential and logarithm functions as sums of infinite matrix products evaluated under $\delta_{1}^{N}$ and $\delta_{1}^{M}$. Pellarin proves somewhat similar looking formulas in the case of the Carlitz module in [22, Thm. 1] and there are some hints towards the formulas of this section contained in the theorems of [16, §6] on trivial MZVs, but in general our formulas are completely original. As discussed in the introduction, we view a special case of these formulas (Corollary 5.10) as giving a function field analogue of the Mellin transform formula for the Riemann zeta function. Thus, we are led to view the map $\delta_{1}^{M}$ as an analogue of integration. This analogy is explored in more detail in Remarks 5.12 and 5.13.
@@
(5.5) $\tilde{\pi}=(-\theta)^{q/(q-1)}\prod_{i=1}^{\infty}\Bigl{(}1-\frac{\theta}{\theta^{q^{i}}}\Bigr{)}^{-1}\in\mathbb{C}_{\infty}$
@@
$p_{n}(\operatorname{Log}_{C}^{\otimes n}(\mathbf{z}))=\delta_{1,\mathbf{z}}^{M}\left(\frac{\tilde{\pi}^{n}}{(t-\theta)\omega_{C}^{n}}\right).$
@@
and also let $u=\frac{\tilde{\pi}^{n}}{t-\theta}$ (by [13, Thm. 6.7] $-u$ is the bottom coordinate of $\mathbf{u}$).
@@
$\delta_{1,\mathbf{z}}^{M}\left(\frac{\tilde{\pi}^{n}}{(t-\theta)\omega_{C}^{n}}\right)=\delta_{1,\mathbf{z}}^{M}\left(\frac{u}{p_{n}(\operatorname{Exp}_{C^{\otimes n}}(\mathbf{u}))}\right)=\Gamma(n)\zeta_{A}(n),$
@@ 6. Log-Algebraic Criterion
$H_{\ell}(x,y)=\sum_{j=0}^{\ell}\sum_{k=1}^{d}\delta_{0}^{N}(\sigma^{-j}(x(h_{k})))\left(\delta_{0}^{M}(\tau^{j-\ell}(y(g_{k})))^{\top}\right)^{(\ell)}.$
@@
$H_{\ell}(x,y;g,h)=\sum_{j=0}^{\ell}\delta_{0}^{N}(\sigma^{-j}(x(h)))\left(\delta_{0}^{M}(\tau^{j-\ell}(y(g)))^{\top}\right)^{(\ell)}.$
@@
$H_{\ell}(bx,y)$ $=\sum_{j=0}^{\ell}\sum_{k=1}^{d}\delta_{0}^{N}(\sigma^{-j}(bx(g_{k})))\left(\delta_{0}^{M}(\tau^{j-\ell}(y(h_{k})))^{\top}\right)^{(\ell)}$
$=\sum_{j=0}^{\ell}\sum_{k=1}^{d}d[b]\delta_{0}^{N}(\sigma^{-j}(x(g_{k})))\left(\delta_{0}^{M}(\tau^{j-\ell}(y(h_{k})))^{\top}\right)^{(\ell)}$
@@
$H_{\ell}(x,by)$ $=\sum_{j=0}^{\ell}\sum_{k=1}^{d}\delta_{0}^{N}(\sigma^{-j}(x(g_{k})))\left(\delta_{0}^{M}(\tau^{j-\ell}(by(h_{k})))^{\top}\right)^{(\ell)}$
$=\sum_{j=0}^{\ell}\sum_{k=1}^{d}\delta_{0}^{N}(\sigma^{-j}(x(g_{k})))\left((d[b]^{\top}\delta_{0}^{M}(\tau^{j-\ell}(y(h_{k}))))^{\top}\right)^{(\ell)}$
$=\sum_{j=0}^{\ell}\sum_{k=1}^{d}\delta_{0}^{N}(\sigma^{-j}(x(g_{k})))\left(\delta_{0}^{M}(\tau^{j-\ell}(y(h_{k})))^{\top}\right)d[b])^{(\ell)}$

### v3

@@ 5. Product Formulas for Exp and Log
In this section we prove formulas which express the exponential and logarithm functions as sums of infinite matrix products evaluated under $\delta_{1}^{N}$ and $\delta_{1}^{M}$. Pellarin proves somewhat similar looking formulas in the case of the Carlitz module in [22, Thm. 1] and there are some hints towards the formulas of this section contained in the theorems of [16, §6] on trivial MZVs, but in general our formulas are completely original. As discussed in the introduction, we view a special case of these formulas (Corollary 5.11) as giving a function field analogue of the Mellin transform formula for the Riemann zeta function. Thus, we are led to view the map $\delta_{1}^{M}$ as an analogue of integration. This analogy is explored in more detail in Remarks 5.13 and 5.14.
@@
Remark 5.9.
We comment that the map $\delta_{1,\mathbf{z}}^{M}$ in the above formula is still the extension described in Definition 2.19; in the second equality one must implicitly use the isomorphism $M\cong\mathbb{C}_{\infty}[t]\cong\mathbb{C}_{\infty}[\tau]^{d}$.
@@
(5.5) $\tilde{\pi}=\theta(-\theta)^{1/(q-1)}\prod_{i=1}^{\infty}\Bigl{(}1-\frac{\theta}{\theta^{q^{i}}}\Bigr{)}^{-1}\in\mathbb{C}_{\infty}$
@@
$p_{n}(\operatorname{Log}_{C}^{\otimes n}(\mathbf{z}))=\delta_{1,\mathbf{z}}^{M}\left((-1)^{n-1}\frac{\tilde{\pi}^{n}}{(\theta-t)\omega_{C}^{n}}\right).$
@@
and also let $u=\frac{\tilde{\pi}^{n}}{\theta-t}$ (by [13, Thm. 6.7] $u$ is the bottom coordinate of $\mathbf{u}$).
@@
$\delta_{1,\mathbf{z}}^{M}\left((-1)^{n-1}\frac{\tilde{\pi}^{n}}{(\theta-t)\omega_{C}^{n}}\right)=\delta_{1,\mathbf{z}}^{M}\left((-1)^{n-1}\frac{u}{p_{n}(\operatorname{Exp}_{C^{\otimes n}}(\mathbf{u}))}\right)=\Gamma(n)\zeta_{A}(n),$
@@ 6. Log-Algebraic Criterion
$H_{\ell}(x,y)=\sum_{j=0}^{\ell}\sum_{k=1}^{d}\delta_{0}^{N}(\sigma^{-j}(y(h_{k})))\left(\delta_{0}^{M}(\tau^{j-\ell}(x(g_{k})))^{\top}\right)^{(\ell)}.$
@@
$H_{\ell}(x,y;g,h)=\sum_{j=0}^{\ell}\delta_{0}^{N}(\sigma^{-j}(y(h)))\left(\delta_{0}^{M}(\tau^{j-\ell}(x(g)))^{\top}\right)^{(\ell)}.$
@@
$H_{\ell}(x,by)$ $=\sum_{j=0}^{\ell}\sum_{k=1}^{d}\delta_{0}^{N}(\sigma^{-j}(by(g_{k})))\left(\delta_{0}^{M}(\tau^{j-\ell}(c(h_{k})))^{\top}\right)^{(\ell)}$
$=\sum_{j=0}^{\ell}\sum_{k=1}^{d}d[b]\delta_{0}^{N}(\sigma^{-j}(y(g_{k})))\left(\delta_{0}^{M}(\tau^{j-\ell}(x(h_{k})))^{\top}\right)^{(\ell)}$
@@
$H_{\ell}(bx,y)$ $=\sum_{j=0}^{\ell}\sum_{k=1}^{d}\delta_{0}^{N}(\sigma^{-j}(y(g_{k})))\left(\delta_{0}^{M}(\tau^{j-\ell}(bx(h_{k})))^{\top}\right)^{(\ell)}$
$=\sum_{j=0}^{\ell}\sum_{k=1}^{d}\delta_{0}^{N}(\sigma^{-j}(y(g_{k})))\left((d[b]^{\top}\delta_{0}^{M}(\tau^{j-\ell}(x(h_{k}))))^{\top}\right)^{(\ell)}$
$=\sum_{j=0}^{\ell}\sum_{k=1}^{d}\delta_{0}^{N}(\sigma^{-j}(y(g_{k})))\left(\delta_{0}^{M}(\tau^{j-\ell}(x(h_{k})))^{\top}\right)d[b])^{(\ell)}$
@@ 6.1. Example: Carlitz Tensor Powers
7. Acknowledgments
The author give sincere thanks to the referee for a very careful reading of the paper and many constructive suggestions. He also thanks Andreas Maurischat for many helpful discussions about the context of the formulas contained in this paper.